OpenStudy (anonymous):

Here is another one, (much easy than the earlier one) How many real \( r \) are there such that the roots of \( x^2 + r x + 6r = 0 \) are both integers?

OpenStudy (mr.math):

By the quadratic formula, \[x={-r\pm \sqrt{r^2-24r} \over 2}\] The roots are integers when \(r^2-24r=(r-12)^2-144\) is a complete square divisible by 4 and r is even.

OpenStudy (mr.math):

perfect square I mean.

OpenStudy (anonymous):

Yes, so how many solutions are there with that constraints?

OpenStudy (mr.math):

There are infinitely many.

OpenStudy (mr.math):

Assuming r is real as indicated in your question.

OpenStudy (anonymous):

There is a finite number of integer solution.

OpenStudy (mr.math):

That's not true. There are infinitely many r's for which \(r(r-24)\) is a perfect squares, but not all of them yield to an integer solution for x.

OpenStudy (mr.math):

I mean what I said was not true @few :D

OpenStudy (mr.math):

24 and 0 are obvious values for r that would give integer solutions.

OpenStudy (asnaseer):

I believe this is just a solution to the Pythagorean triples such that:  and the only such triples are: 12,12,0 => r-12=12 5,12,13 => r-12=13 9,12,15 => r-12=15 12,16,20 => r-12=20 12,35,37 => r-12=37 giving solutions of: r=24,25,27,32,49 plus r=0 is a trivial solution

OpenStudy (mr.math):

This is very nice, but someone has to be familiar with these numbers. Plus we need not to forget about negative values of r.

OpenStudy (mr.math):

For example r=-3 is a solution.

OpenStudy (mr.math):

So is r=-1.

OpenStudy (asnaseer):

true, negative r would give us: |r-12|=12, 13, 15, 20, 37 giving us the following additional solutions: r=-1, -3, -8, -25

OpenStudy (mr.math):

Yeah! I should have seen it :D

OpenStudy (anonymous):

@asnaseer:Magnum opus! Indeed Pythagorean triples are the only easy way out.

OpenStudy (asnaseer):

thx - you're teaching me Latin now :-D

OpenStudy (anonymous):

Auf Wiedersehen ;)